OpenStudy (anonymous):

PLEASE HELP!! A rain gutter is to be constructed of aluminum sheets 12 inches wide. After marking off a length of 4 inches from each edge, this length is bent up at an angle Θ. The area A of the opening may be expressed as the function: A(Θ) = 16 sin Θ • (cos Θ + 1). If Θ = 90°, what is the area of the opening? When i plug 90 in i keep getting 15, but my choices are: 21.8 in2 16.0 in2 7.9 in2 1.0 in2

OpenStudy (anonymous):

@dangerousjesse any luck?

OpenStudy (anonymous):

@phi @vineeth10

OpenStudy (anonymous):

@ganeshie8 @MATHMALE

OpenStudy (phi):

first, you can do this without a calculator. But if you do use a calculator, be sure it is in degree mode. you should(!) know sin(90º)= 1 and cos(90º)= 0

OpenStudy (anonymous):

@phi i just figured it out actually!! C right? and could you please help me with one more!! im stuck on it :/

OpenStudy (anonymous):

Simplify \[\sqrt{\left( 1-\sin \Theta \right)\left( 1+\sin \Theta \right)}\]

OpenStudy (anonymous):

@phi i have no idea what to do

OpenStudy (phi):

First, How did you get C 7.9? It seems if sin(90) is 1 and cos(90) is 0 you would do 16 sin Θ • (cos Θ + 1) 16 • 1 •(0+1) and now we add 0+1 (not *too* hard) and multiply that by 16*1 you don't get 7.9

OpenStudy (anonymous):

For some reason i divided 16 in 2 idk why haha thank you!!

OpenStudy (phi):

any idea what the answer is for your first question?

OpenStudy (anonymous):

b right?

OpenStudy (anonymous):

any idea on how to solve the second one? i really have no idea what to do @phi

OpenStudy (phi):

do you know how to multiply the two binomials inside the square root (some people use FOIL)

OpenStudy (anonymous):

yeah when i foil i get sqrt(1-sin O)

OpenStudy (anonymous):

is that right? @phi

OpenStudy (phi):

can you show your work?

OpenStudy (anonymous):

F: 1*1=1 O: 1*sinO=sinO I: -sinO*1=-sinO L: -sinO*sinO=-sin2 so its sqrt(1-sin^2) oops!

OpenStudy (anonymous):

these are my options btw @phi ±sin Θ |cos Θ| ±tan Θ square root sine theta

OpenStudy (phi):

ok looks better there is a *very useful* identity you should remember for trig problems \[ \sin^2 x + \cos^2 x = 1 \] which you can rewrite as \[ \sin^2 x = 1 - \cos^2 x \] or \[ \cos^2 x= 1 - \sin^2 x \]

OpenStudy (anonymous):

would it be B?

OpenStudy (phi):

yes. replace 1 -sin^2 with cos^2 inside the square root sqrt(cos^2) = cos is the only choice that looks close (they must only want the positive root) | cos Θ |

OpenStudy (anonymous):

@phi thank you so so much!! i really truly appreciate it :)

OpenStudy (phi):

yw